Consider the configuration shown below, with a negative charge Q0 at position x0=0 and another positive charge Q1 at position 1 along the x-axis. Determine the position At which position x2 could another positive charge charge Q2 be added, such that the total force on Q0 is 0? Give two examples of x2 and corresponding Q2, in terms of x0, x1, Q0 and Q1 (or a subset of these variables). 


1
Expert's answer
2022-03-24T16:36:55-0400

Answer

Ratio of charge is

"\frac{Q_1}{Q_2}=\frac{x_0^2}{x_1^2}"

So electrostatic force is

"F=\frac{kQQ'}{r^2}"